OpenStudy (anonymous):

solve equation by square root property (x - 9)sqrt 2 = 49

OpenStudy (anonymous):

@jim_thompson5910

geerky42 (geerky42):

Like that? \[(x-9)\sqrt2 = 49\]

geerky42 (geerky42):

or do you mean \(\huge \sqrt[x-9]{2} = 49\) ?

OpenStudy (anonymous):

no like (x - 9) and then a 2 up top and then = 49

geerky42 (geerky42):

\(\Large (x-9)^{\sqrt{2}} = 49\)?

OpenStudy (anonymous):

just a regular 2

OpenStudy (anonymous):

but if that means sq rt 2 then ya lol

geerky42 (geerky42):

Ok so \((x-9)^2 = 49\). Take square root of both sides. \(x - 9 = \pm7\) \(x = 9 \pm7\) \(x = 9 + 7\) or \(9-7\) \(x = 16 \) or \(2\)